# Low-Rank and Sparse Model Merging for Multi-Lingual Speech Recognition and Translation

### v2

@@ Abstract
Language diversity presents a significant challenge in speech-to-text (S2T) tasks, such as automatic speech recognition and translation. Traditional multi-task training approaches aim to address this by jointly optimizing multiple speech recognition and translation tasks across various languages. While models like Whisper, built on these strategies, demonstrate strong performance, they still face issues of high computational cost, language interference, suboptimal training configurations, and limited extensibility. To overcome these challenges, we introduce LoRS-Merging (low-rank and sparse model merging), a novel technique designed to efficiently integrate models trained on different languages or tasks while preserving performance and reducing computational overhead. LoRS-Merging combines low-rank and sparse pruning to retain essential structures while eliminating redundant parameters, mitigating language and task interference, and enhancing extensibility. Experimental results across a range of languages demonstrate that LoRS-Merging reduces the word error rate by 10% and improves BLEU scores by 4% compared to conventional multi-lingual multi-task training baselines. Our findings suggest that model merging, particularly LoRS-Merging, is a scalable and effective complement to traditional multi-lingual training strategies for S2T applications111The detailed data and code will be released at [URL].
@@
Qiuming Zhao1 Guangzhi Sun2 Chao Zhang1 Mingxing Xu1 Thomas Fang Zheng1††thanks: Correspondence 1Tsinghua University, China 2University of Cambridge, United Kingdom zqm23@mails.tsinghua.edu.cn, gs534@cam.ac.uk, {cz277,xumx,fzheng}@tsinghua.edu.cn
@@ 1 Introduction
Language diversity poses a significant challenge in speech-to-text (S2T) tasks, such as automatic speech recognition (ASR) Prabhavalkar et al. (2023) and speech translation (ST) Xu et al. (2023). With over 7,000 languages spoken worldwide, developing robust S2T systems that generalise across varied linguistic structures remains a fundamental research goal Liu and Niehues (2024); Cheng et al. (2023); Sun et al. (2023); Saif et al. (2024); Wang et al. (2021); Le et al. (2021). The advent of end-to-end (E2E) models Chan et al. (2016); Gulati et al. (2020); Barrault et al. (2023) has marked a paradigm shift in S2T tasks, enabling direct mapping from speech to text across multiple languages within a unified framework. A prominent example is Whisper Radford et al. (2023), an advanced multi-lingual speech model trained on a large-scale, diverse dataset covering multiple languages and tasks. Despite these advances, existing multi-lingual models still encounter significant challenges in scalability, efficiency, and performance trade-offs.
@@
To address these challenges, multi-lingual training strategies Saif et al. (2024); Xiao et al. (2021); Bai et al. (2018) have been adopted, aiming to enhance model generalisation across languages. These approaches typically rely on joint optimisation of diverse S2T tasks across multiple languages, leveraging shared representations to improve performance. Nevertheless, multi-lingual training is subject to inherent limitations, including substantial training costs, complex model configurations, and limited access to training data across multiple languages and tasks. Moreover, when handling new languages, the training methods typically require training from scratch.
@@
To mitigate these issues, this paper proposes to use model merging Ilharco et al. (2023); Yang et al. (2024a); Khan et al. (2024) to integrate models trained on different languages or tasks while maintaining performance and reducing computational overhead. Model merging merges the parameters of multiple separate models with different capabilities to build a universal model. With its high flexibility, model merging enables the seamless incorporation of new languages or tasks without the need for retraining the entire model. Additionally, since model merging allows models for different languages or tasks to be trained independently, it can effectively alleviate negative transfer issues Wang et al. (2019); Zhang et al. (2023b); Wang et al. (2020b) commonly observed in multi-lingual training. This training independence also enables the use of optimal training configurations for each language or task instead of the unified settings required in multi-lingual training.
@@
Moreover, we propose Low-Rank and Sparse model Merging (LoRS-Merging), which uses a low-rank component to capture the compact structure and a sparse component to capture the scattered details in the weights. LoRS-Merging retains effective parts of structure and details while reducing redundant parts to reduce task interference. Specifically, coarse-grained singular value pruning is used to retain the low-rank structure, while fine-grained magnitude pruning is used to remove redundant details. The main contribution of this paper can be summarised as follows.
We propose LoRS-Merging, a low-rank and sparse model merging method for multi-lingual ASR and speech translation. To the best of our knowledge, LoRS-Merging is the first work that explores model merging for speech models.
@@
LoRS-Merging exploits the combination of low-rank structure and sparsity of language-specific and task-specific weights in model merging, minimising the parameter redundancy and conflicts as well as providing an efficient way to incorporate new knowledge from a task or language-specialised model.
Experiments are performed across 10 different languages where LoRS-Merging achieves 10% relative WER reduction and 4% relative BLEU increase compared to the multi-lingual multi-task training baseline. Moreover, we show that negative interference largely exists in multi-lingual training and LoRS-Merging alleviates this issue.
@@ 2.1 Multi-Lingual ASR and ST
The more effective approach is to adopt multi-lingual training strategies, such as multi-objective optimisation Saif et al. (2024); Zhang et al. (2022), adversarial learning Xiao et al. (2021), meta learning Hsu et al. (2020), and reinforcement learning Bai et al. (2018). Moreover, large-scale pre-training by leveraging massive amounts of multi-lingual and multi-task data enables models to learn robust and transferable representations across languages, e.g. Whisper Radford et al. (2023), SeamlessM4T Barrault et al. (2023), and AudioPaLM Rubenstein et al. (2023). LoRS-Merging, as an efficient post-training method proposed in this paper, further advances multi-lingual ASR and ST based on pre-trained speech models.
@@ 2.2 Model Merging
Another stream considers the model spaces, including activation spaces and weight spaces. Research on activation spaces seeks to align the output representations or loss of the merged model with those of each single model as closely as possible Yang et al. (2024b); Wei et al. (2025); Xiong et al. (2024). Studies based on weight spaces aim to remove redundant parameters or localise effective parameters to resolve task interference. TIES-Merging Yadav et al. (2024) and DARE Yu et al. (2024) perform magnitude or random pruning on each single model to significantly remove redundant parameters. TALL-masks Wang et al. (2024) and Localise-and-Stitch He et al. (2024) optimise binary masks to localise sparse and effective task-specific parameters. In contrast, LoRS-Merging explores weight space merging by considering not only the detailed parameter redundancy as well as maintaining the effective structure of the weight space via low-rank pruning.
@@ 3.2 Model Merging for Speech Models
The model merging process for speech model on S2T tasks with LoRS-Merging as an example is shown in Fig. 1, which comprises four steps. In step 1, a suitable pre-trained speech model is selected. In step 2, for each target language and target task combination, e.g. Catalan ASR, the pre-trained model is finetuned with the task-language-specific data and the delta weight is obtained. In step 3, weight pruning is applied to remove redundant and conflicting delta parameters. In step 4, task arithmetic is applied to combine pruned delta weights into each single merged matrix and hence obtain the merged model.
Model merging allows new language or task knowledge to be integrated into the model in a flexible post-training manner. When a new set of data for a specific language is obtained, model merging incorporates such knowledge by fine-tuning with the new data alone with data-specific configuration, which also releases the burden of requiring other data to avoid catastrophic forgetting. This benefit is thoroughly demonstrated in our experiments.
@@ 3.3 Low-Rank and Sparse Model Merging
Compared to multi-lingual or multi-task training methods, model merging is a simpler and more efficient approach, enabling the seamless incorporation of new languages or tasks without the need for retraining. Additionally, due to its training independence, it mitigates negative transfer and provides optimal training configurations for each language or task to improve performance.
@@ 4.1 Data
CoVoST-2 Wang et al. (2020a) is a large-scale multi-lingual ST corpus based on Common Voice. It covers translations from English into 15 languages and from 21 languages into English, with a total of 2,880 hours of speech from 78k speakers. We selected 5 high-resource languages and 5 low-resource languages as two language sets to investigate their ASR tasks and the from X to English ST tasks. The high-resource language set includes Catalan (ca), German (de), Spanish (es), French (fr), and Italian (it), while the low-resource language set includes Indonesian (id), Dutch (nl), Portuguese (pt), Russian (ru), and Swedish (sv). Due to the more abundant data in the high-resource language set, our main experimental results are obtained on the high-resource language set, while the low-resource language set serves as an auxiliary validation set. To balance the amount of data across different languages, we fixed the duration of traning data for each language, with 5 hours for the high-resource language set and 1 hour for the low-resource language set. The dev and test sets of both language sets are 1 hour in duration.
@@ 4.2 Model and Training Specifications
For each language-specific or task-specific finetuned model, we use a different, optimal learning rate for each during training, and these models are subsequently used for model merging. Finetuning involves updating all parameters. We choose Adam as the optimiser, set the batch size to 8, the accumulation iterations to 4, and train for 10 epochs. We also select the proportions of low-rank parameters retained by SVP from {1%, 2%, 3%, 5%} and sparse parameters retained by MP from {10%, 20%, 40%, 60%}. The beam size for decoding is set to 20 across all languages and tasks. We use Sclite and SacreBLEU tools to score the ASR and ST results, respectively. See Appendix A for more details on hyper-parameter settings. Our experiments are performed on a single RTX 4090 GPU where training on one language and one task with 5 hours of speech data requires 1 hour.
@@ 4.3 Baseline and Merging Methods
We use Multi-lingual and multi-task training as the baseline for comparison with model merging methods, where training is conducted on data mixed from both multi-lingual and multi-task sets. To ensure a fair comparison, the same amount of training data is used from each language and each task. Note that for 5 different languages with both ASR and ST tasks, multi-lingual and multi-task training is performed on 10 times more data and hence 10 times more computational resources.
@@
Weight Averaging (WA) merges multiple single models by and unweighted averaging of their weights, .
Task Arithmetic (TA) uses a scaling factor to weight multiple task vectors estimated on a small development set, .
MP-Merging performs fine-grained magnitude pruning on task vectors to reduce redundancy at the detail level, .
SVP-Merging performs coarse-grained singular value pruning on task vectors to reduce redundancy at the structure level, . (see Section 3.1.2).
Moreover, we compare methods against the performance of fine-tuning on each language-task combination. This is the topline of all merging methods since the model is completely adapted to a specific language for a specific task with optimised configurations and without any language conflicts.
@@ 5.1 Multi-Lingual Model Merging
System WER ca de es fr it Avg. Pretrained 20.6 19.6 14.7 24.5 19.4 19.88 Finetuned 19.5 19.7 14.4 22.1 19.2 19.05 Multi-lingual training 17.1 21.8 15.1 22.6 21.9 19.69 Weight Averaging 19.1 19.1 14.2 24.5 20.3 19.55 Task Arithmetic 19.1 18.8 13.9 24.0 19.8 19.23 MP-Merging 19.4 19.3 14.0 23.8 18.1 19.03 SVP-Merging 19.5 19.3 14.2 23.6 18.4 19.11 LoRS-Merging 19.0 18.8 13.9 23.5 18.5 18.85
System BLEU ca de es fr it Avg. Pretrained 21.1 24.1 28.6 26.8 26.8 25.48 Finetuned 22.6 24.6 29.2 27.2 27.3 26.18 Multi-lingual training 21.4 24.4 28.8 26.8 27.2 25.72 Weight Averaging 22.3 24.1 28.6 27.2 26.9 25.82 Task Arithmetic 22.1 24.3 28.9 27.3 26.8 25.88 MP-Merging 22.1 24.7 28.9 27.3 26.9 25.98 SVP-Merging 22.1 24.7 29.0 27.4 26.8 26.00 LoRS-Merging 22.2 24.8 29.0 27.5 26.9 26.08
@@
Language knowledge interference yields imbalanced improvements: Table 1 shows the multi-lingual results of the ASR task with the high-resource language set. On average, multi-lingual training slightly improves the pretrained model but significantly underperforms the finetuned models and merging methods. This may be due to negative interference between the knowledge of different languages, leading to gradient conflicts during training Wang et al. (2020b). From a per-language perspective, it is observed that ca and fr achieve the largest improvements during fine-tuning while still showing significant improvements in multi-lingual training, whereas languages with smaller improvements during finetuning exhibit a substantial performance drop in multi-lingual training, even worse than the pretrained model. This indicates a strong language conflict in multi-lingual training, with ca and fr dominating. Additionally, we observe that the optimal learning rates for finetuned models vary significantly across languages (see Appendix A), while the unified learning rate configuration required by multi-lingual training prevents each language from reaching its optimal performance.
Model merging mitigates language conflicts: In contrast, model merging methods show obvious improvements across almost all languages, demonstrating reduced conflict and better stability. Among model merging methods, TA outperforms WA due to its flexible scaling factor. Both MP-Merging and SVP-Merging further improve the performance of TA by reducing redundancy, and MP-Merging slightly outperforms SVP-Merging due to its finer-grained pruning. Combining the advantages of SVP and MP, LoRS-Merging achieves the best performance.
Table 2 provides the multi-lingual results on ST task with the high-resource language set. The main conclusion is consistent with the ASR task: model merging methods still significantly outperform multi-lingual training, with LoRS-Merging achieving the best performance.

### v3

@@ Abstract
Language diversity presents a significant challenge in speech-to-text (S2T) tasks, such as automatic speech recognition and translation. Traditional multi-lingual multi-task training approaches aim to address this by jointly optimising multiple speech recognition and translation tasks across various languages. While models like Whisper, built on these strategies, demonstrate strong performance, they still face issues of high computational cost, language interference, suboptimal training configurations, and limited extensibility. To overcome these challenges, we introduce LoRS-Merging (low-rank and sparse model merging), a novel technique designed to efficiently integrate models trained on different languages or tasks while preserving performance and reducing computational overhead. LoRS-Merging combines low-rank and sparse pruning to retain essential structures while eliminating redundant parameters, mitigating language interference, and enhancing extensibility. Experimental results across 10 languages demonstrate that LoRS-Merging significantly outperforms multi-lingual multi-task training, sequential training, and other merging methods, achieving over 20% improvement in normalised performance. Our findings suggest that model merging, particularly LoRS-Merging, is a scalable and effective complement to traditional multi-lingual training strategies for S2T applications111The detailed data and code will be released at [URL].
@@
Qiuming Zhao1 Guangzhi Sun2 Chao Zhang1††thanks: Correspondence 1Tsinghua University, China 2University of Cambridge, United Kingdom zqm23@mails.tsinghua.edu.cn, gs534@cam.ac.uk, cz277@tsinghua.edu.cn
@@ 1 Introduction
Language diversity poses a significant challenge in speech-to-text (S2T) tasks, such as automatic speech recognition (ASR) Prabhavalkar et al. (2023) and speech translation (ST) Xu et al. (2023). With over 7,000 languages spoken worldwide, developing robust S2T systems that generalise across varied linguistic structures remains a fundamental research goal Liu and Niehues (2024); Cheng et al. (2023); Sun et al. (2023); Saif et al. (2024); Wang et al. (2021); Le et al. (2021). The transition from pipeline systems to end-to-end (E2E) models Chan et al. (2016); Gulati et al. (2020); Barrault et al. (2023) has marked a paradigm shift in S2T tasks, enabling direct speech-to-text mapping across multiple languages within a unified framework. A prominent example is Whisper Radford et al. (2023), an advanced multi-lingual speech model trained on a large-scale, diverse dataset covering multiple languages and tasks. Despite these advances, existing multi-lingual models still encounter significant challenges in scalability, efficiency, and performance trade-offs.
@@
To address these challenges, multi-lingual training strategies Saif et al. (2024); Xiao et al. (2021); Bai et al. (2018) have been adopted, aiming to enhance model generalisation across languages. These approaches typically rely on joint optimisation of diverse S2T tasks across multiple languages, leveraging shared representations to improve performance. Nevertheless, multi-lingual training is subject to inherent limitations, including substantial training costs, complex model configurations, and limited access to training data across multiple languages and tasks. Moreover, when handling new languages, the training methods typically require training from scratch, leading to poor extensibility.
@@
To mitigate these issues, this paper proposes to use model merging Ilharco et al. (2023); Yang et al. (2024a); Khan et al. (2024) to integrate models trained on different languages or tasks while maintaining performance and reducing computational overhead. Model merging merges the parameters of multiple separate models with different capabilities to build a universal model. With its high flexibility, model merging enables the seamless incorporation of new languages or tasks without the need for retraining the entire model. Additionally, since model merging allows models for different languages or tasks to be trained independently, it can effectively alleviate negative transfer issues Wang et al. (2019); Zhang et al. (2023b); Wang et al. (2020b) commonly observed in multi-lingual training. This training independence also enables optimal training configurations for each language or task to improve performance, instead of the unified settings required in multi-lingual training.
@@
Moreover, we propose Low-Rank and Sparse model Merging (LoRS-Merging), which uses a low-rank component to capture the compact structure and a sparse component to capture the scattered details in the weights. LoRS-Merging retains effective parts of structure and details while reducing redundant parts to reduce language interference. Specifically, coarse-grained singular value pruning is used to retain the low-rank structure, while fine-grained magnitude pruning is used to remove redundant details. The main contribution of this paper can be summarised as follows.
To the best of our knowledge, this work is the first to explore model merging for speech-to-text models. Specifically, we treat speech tasks (recognition and translation) and different languages as two separate merging levels and explore different hierarchies for model merging.
@@
We propose LoRS-Merging, which exploits the low-rank structure and sparsity of model weights to minimise model redundancy and language conflicts as well as providing an efficient way to incorporate new knowledge from a task- or language-specialised model.
Experiments are performed across 10 languages, where LoRS-Merging significantly outperforms multi-lingual multi-task training, sequential training, and other merging methods, achieving over 20% improvement in normalised performance.
@@ 2.1 Multi-Lingual ASR and ST
The more effective approach is to adopt multi-lingual training strategies, such as multi-objective optimisation Saif et al. (2024); Zhang et al. (2022), adversarial learning Xiao et al. (2021), meta learning Hsu et al. (2020), and reinforcement learning Bai et al. (2018). Moreover, large-scale pretraining by leveraging massive amounts of multi-lingual and multi-task data enables models to learn robust and transferable representations across languages, e.g. Whisper Radford et al. (2023), SeamlessM4T Barrault et al. (2023), and AudioPaLM Rubenstein et al. (2023). LoRS-Merging, as an efficient post-training method proposed in this paper, further advances multi-lingual ASR and ST based on pretrained speech models.
@@ 2.2 Model Merging
Another stream considers the model spaces, including activation spaces and weight spaces. Research on activation spaces seeks to align the output representations or loss of the merged model with those of each single model as closely as possible Yang et al. (2024b); Wei et al. (2025); Xiong et al. (2024). Studies based on weight spaces aim to localise effective parameters or remove redundant parameters to resolve task interference. TALL-masks Wang et al. (2024) and Localise-and-Stitch He et al. (2024) optimise binary masks to localise sparse and effective task-specific parameters. TIES-Merging Yadav et al. (2024) and DARE Yu et al. (2024) perform magnitude or random pruning on each single model to reduce redundancy at the detailed parameter level. TSV-M Gargiulo et al. (2024), on the other hand, adopts singular value pruning to reduce redundancy at the structural level. In contrast, LoRS-Merging explores weight space merging by considering not only the detailed parameter redundancy as well as maintaining the effective structure of the weight space.
@@ 3.2 Model Merging for Speech Models
The model merging process for speech model on S2T tasks with LoRS-Merging as an example is shown in Fig. 1, which comprises four steps. In step 1, a suitable pretrained speech model is selected. In step 2, for each target language and target task combination, e.g. Catalan ASR, the pretrained model is finetuned with the task-language-specific data and the delta weight is obtained. In step 3, weight pruning is applied to remove redundant and conflicting delta parameters. In step 4, task arithmetic is applied to combine pruned delta weights into each single merged matrix and hence obtain the merged model.
Model merging allows new language or task knowledge to be integrated into the model in a flexible post-training manner. Compared to multi-lingual or multi-task training methods, model merging is a simpler and more efficient approach, enabling the seamless incorporation of new languages or tasks without the need for retraining. Additionally, due to its training independence, it mitigates language conflicts and provides optimal training configurations for each language or task to improve performance. Compared to sequential training, model merging eliminates the need for additional training data to avoid catastrophic forgetting. Our experiments thoroughly demonstrate these benefits.
@@ 4.1 Data
CoVoST-2 Wang et al. (2020a) is a large-scale multi-lingual ST corpus based on Common Voice. It covers translations from English into 15 languages and from 21 languages into English, with a total of 2,880 hours of speech from 78k speakers. We selected 5 high-resource languages and 5 low-resource languages as two language sets to investigate their ASR tasks and the from X to English ST tasks. The high-resource language set includes Catalan (ca), German (de), Spanish (es), French (fr), and Italian (it), while the low-resource language set includes Indonesian (id), Dutch (nl), Portuguese (pt), Russian (ru), and Swedish (sv). Due to the more abundant data in the high-resource language set, our main experimental results are obtained on the high-resource language set, while the low-resource language set serves as an auxiliary evaluation set. To balance the amount of data across different languages, we fixed the duration of traning data for each language, with 5 hours for the high-resource language set and 1 hour for the low-resource language set. The dev and test sets of both language sets are 1 hour in duration.
@@ 4.2 Model and Training Specifications
For each language-specific or task-specific finetuned model, we use a different, optimal learning rate for each during training, and these models are subsequently used for model merging. Finetuning involves updating all parameters. We choose Adam as the optimiser, set the batch size to 8, the accumulation iterations to 4, and train for 10 epochs. The beam size for decoding is set to 20 across all languages and tasks. We use Sclite and SacreBLEU tools to score the ASR and ST results, respectively. In addition, we perform statistical significance testing using a paired bootstrap test with 1,000 resampling iterations, each sampling 300 examples with replacement, and report the results in the caption of each table. See Appendix A for more details on the experimental setup. Our experiments are performed on a single RTX 4090 GPU where training on one language and one task with 5 hours of speech data requires 1 hour.
@@ 4.3 Baseline and Merging Methods
We use the pretrained model as the baseline and multi-lingual multi-task training as the stronger baseline, which is trained on data mixed from both multi-lingual and multi-task sets. Note that finetuned models are typically available, so model merging requires no finetuning and only adjusts merging coefficients on a small development set. Even if finetuning is needed, the complex training configurations of multi-lingual and multi-task training require more hyperparameter tuning steps. Overall, model merging consumes significantly fewer computational resources than multi-lingual and multi-task training.
@@
Weight Averaging (WA) Wortsman et al. (2022) merges multiple single models by a uniform averaging of their weights.
Task Arithmetic (TA) Ilharco et al. (2023) uses a scaling factor, estimated on a small development set, to weight multiple task vectors.
TIES-Merging Yadav et al. (2024) merges single models via Trim, Elect, and Disjoint Merge steps to reduce parameter interference.
DARE Yu et al. (2024) drops and rescales delta parameters to mitigate parameter interference.
TSV-M Gargiulo et al. (2024) proposes Task Singular Vectors and reduces structural redundancy to reduce task interference.
In addition, we report the normalised performance difference defined in Eqn. (9).
where is the performance metric of the target system, and are for the pretrained (baseline performance) and finetuned (topline performance) systems respectively. Note that better reflects the performance gains for model merging since the pretrained system already achieves competitive performance.
@@ 5.1 Multi-Lingual Model Merging
System WER ca de es fr it Avg. Pretrained 20.6 19.6 14.7 24.5 19.4 19.88 - Finetuned 19.5 19.7 14.4 22.1 19.2 19.05 100.0% Multi-lingual training 17.1 21.8 15.1 22.6 21.9 19.69 22.9% Sequential training 20.6 19.6 14.6 24.4 19.4 19.84 4.8% Weight Averaging 19.1 19.1 14.2 24.5 20.3 19.55 39.8% Task Arithmetic 19.1 18.8 13.9 24.0 19.8 19.23 78.3% TIES-Merging 19.3 19.3 13.9 23.8 18.1 18.99 107.2% DARE 18.9 18.9 13.9 23.8 19.8 19.16 86.7% TSV-M 19.5 19.5 14.1 23.5 18.4 19.10 94.0% LoRS-Merging 18.9 18.8 13.9 23.6 18.1 18.77 133.7%
System BLEU ca de es fr it Avg. Pretrained 21.1 24.1 28.6 26.8 26.8 25.48 - Finetuned 22.6 24.6 29.2 27.2 27.3 26.18 100.0% Multi-lingual training 21.4 24.4 28.8 26.8 27.2 25.72 34.3% Sequential training 21.5 24.3 28.9 26.9 27.3 25.78 42.9% Weight Averaging 22.3 24.1 28.6 27.2 26.9 25.82 48.6% Task Arithmetic 22.1 24.3 28.9 27.3 26.8 25.88 57.1% TIES-Merging 22.1 24.7 29.0 27.1 26.9 25.96 68.6% DARE 22.1 24.5 28.9 27.2 26.8 25.90 60.0% TSV-M 22.0 24.6 29.0 27.3 26.8 25.94 65.7% LoRS-Merging 22.4 24.8 28.9 27.6 27.0 26.14 94.3%
@@
Language knowledge interference yields imbalanced improvements: Table 1 shows the multi-lingual results of the ASR task with the high-resource language set. On average, multi-lingual training slightly improves the pretrained model but significantly underperforms the finetuned models and merging methods. This may be due to negative interference between the knowledge of different languages, leading to gradient conflicts during training Wang et al. (2020b). From a per-language perspective, it is observed that ca and fr achieve the largest improvements during finetuning while still showing significant improvements in multi-lingual training, whereas languages with smaller improvements during finetuning exhibit a substantial performance drop in multi-lingual training, even worse than the pretrained model. This indicates a strong language conflict in multi-lingual training, with ca and fr dominating. Additionally, we observe that the optimal learning rates for finetuned models vary significantly across languages (see Appendix A), while the unified learning rate configuration required by multi-lingual training prevents each language from reaching its optimal performance. Moreover, the substantially inferior performance of sequential training indicates the presence of catastrophic forgetting.
Model merging mitigates language conflicts: In contrast, model merging methods show significant improvements across almost all languages, demonstrating reduced conflict and better stability. Among model merging methods, TA outperforms WA due to its flexible scaling factor. TIES-Merging and DARE reduce redundancy at the detail level, while TSV-M addresses redundancy at the structural level, thereby achieving obvious improvements over TA. Furthermore, LoRS-Merging reduces redundancy at the detail level while preserving critical structures, resulting in the best performance.
Table 2 provides the multi-lingual results on ST task with the high-resource language set. The main conclusion is consistent with the ASR task: model merging methods still significantly outperform multi-lingual training and sequential training, with LoRS-Merging achieving the best performance, demonstrating superior multi-lingual and continual learning capabilities.